| x| note x^2 means x square

| x| <1;let y =1+x+x^2+…………to infinity then dy/dx equal to


note x^2 means   x square

Grade:

1 Answers

AskIITian Expert Priyasheel - IITD
14 years ago

1+x+x^2+... =1/(1-x). (It is an infinite geometric series and given mod(x) < 1)

y=1/(1-x),

dy/dx = {(1-x)0 - 1(-1)}/(1-x)^2

         = 1/(1-x)^2